Calculate the pH for each case in the titration of 50.0 mL of 0.230 M HClO(aq)0.230 M HClO(aq) with 0.230 M KOH(aq).0.230 M KOH(aq). Use the ionization constant 4.0×10–8 for HClO.

What is the pH after the addition of 50.0 mL KOH?
Answer #1

Given:
M(HClO) = 0.23 M
V(HClO) = 50 mL
M(KOH) = 0.23 M
V(KOH) = 50 mL


mol(HClO) = M(HClO) * V(HClO)
mol(HClO) = 0.23 M * 50 mL = 11.5 mmol

mol(KOH) = M(KOH) * V(KOH)
mol(KOH) = 0.23 M * 50 mL = 11.5 mmol


We have:
mol(HClO) = 11.5 mmol
mol(KOH) = 11.5 mmol

11.5 mmol of both will react to form ClO- and H2O

ClO- here is strong base
ClO- formed = 11.5 mmol
Volume of Solution = 50 + 50 = 100 mL
Kb of ClO- = Kw/Ka = 1*10^-14/4*10^-8 = 2.5*10^-7
concentration ofClO-,c = 11.5 mmol/100 mL = 0.115M

ClO- dissociates as

ClO-        + H2O   ----->     HClO +   OH-
0.115                        0         0
0.115-x                      x         x


Kb = [HClO][OH-]/[ClO-]
Kb = x*x/(c-x)
Assuming x can be ignored as compared to c
So, above expression becomes

Kb = x*x/(c)
so, x = sqrt (Kb*c)
x = sqrt ((2.5*10^-7)*0.115) = 1.696*10^-4

since c is much greater than x, our assumption is correct
so, x = 1.696*10^-4 M



[OH-] = x = 1.696*10^-4 M

use:
pOH = -log [OH-]
= -log (1.696*10^-4)
= 3.7707


use:
PH = 14 - pOH
= 14 - 3.7707
= 10.2293
Answer: 10.23